Previous

Calendar

February

2015

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

1

JD32

2

JD33

3

JD34
Site Visit (JK): Details

JK added new extention cord from tower to the picarro shed

4

JD35

5

JD36
Site Visit (ES): Details

ES updated EC155 firmware and for 27m-level CSATs: switched CSAT3B embedded code from v5.0 to v4.0 and CSAT3 #3 from v4.0 to v5.0

6

JD37

7

JD38
Site Visit (SpB): Details , Photos (32)

Span cylinder at 250 psi. N2 at 600 psi

moved cc319297 (n2) to tower (along fence)

re-started the picarro system (moved from DB shed)

at around 14:00 MST, re-started LI-7500A OP1

from approx 15:01-15:15 MST, Changed 2-micron Nupro filter in profile (li-6251) box

from 15:30-15:40 MST, swiped krypton hygrometer

CU Prius

very windy, very warm (clear skies)

8

JD39
very windy

9

JD40

10

JD41

11

JD42

12

JD43
Site Visit (ES): Details

ES updated tubing/plumbing in EC155 (note, the CSAT3A boom might have been moved)

13

JD44
Site Visit (MR): Details

MR downloaded data/photos

14

JD45

15

JD46

16

JD47

17

JD48

18

JD49

19

JD50
Chinese New Year (Wood Sheep)

20

JD51
Site Visit (SpB): Details , Photos (23)

Span cylinder at 40 --> 1800 psi. N2 at 250 --> 1850 psi

at approx 10:30 MST, removed cc316422 (n2) started cc197414 at 1850 psi

at approx 10:30 MST, removed cc254814 (span) started cc185237 (dave bowling cylincer) at 1800 psi

at around 10:00 MST, disconnected power from LI-7500A OP1

swapped SD card from tower-base cam

SpB Jeep

clear in early AM, then heavy snow

21

JD52

22

JD53

23

JD54
SpB Off

24

JD55

25

JD56

26

JD57
significant snow

27

JD58

28

JD59
at 21:27 MST, started ops75: changed EOL CSAT3s to 60-Hz sample rate